OpenStudy (anonymous):

How would you go about working a proportion, such as 6/2=4/p?

OpenStudy (texaschic101):

6/2 = 4/p (cross multiply) (6)(p) = (2)(4) solve for p....can you take it from here ?

OpenStudy (anonymous):

So you would have 6p=8, and then would you divide by 6?

OpenStudy (texaschic101):

yes

OpenStudy (anonymous):

so p would equal 1.3?

OpenStudy (raden_zaikaria):

better put 4/3

OpenStudy (raden_zaikaria):

p = 4 over 3

OpenStudy (anonymous):

Where did the 3 come from?

OpenStudy (texaschic101):

6p = 8 p = 8/6 = 4/3 check... 6/2 = 4 / (4/3) 3 = 4 * 3/4 3 = 12/4 3 = 3 (correct) answer : p = 4/3

OpenStudy (anonymous):

Ohhhh. Okay, now I get where I when wrong. Thank you. :)

OpenStudy (texaschic101):

your welcome....I am glad you understand